# Stars and quark stars in bumblebee gravity

### v2

@@ Abstract
In this work, the interior spacetime of stars is built in a Lorentz symmetry breaking model called bumblebee gravity. Firstly, we calculated the modified Tolman-Oppenheimer-Volkoff equation in this context of modified gravity. Then we show that the bumblebee field, responsible for the symmetry breaking, increases the star mass-radius relation when it assumes its vacuum expectation value. When compared to the general relativity mass-radius relation, a Lorentz symmetry breaking context, like the bumblebee gravity, could provide more massive stars, surpassing the $2.5M_{\odot}$ limit as the interior of the star is described by quark matter with the MIT bag model.
@@ I Introduction
In this sense, quark matter in astrophysics has been subject of many papers. In modified theories of gravity, we see works in Einstein-Gauss-Bonnet gravity Gammon:2023uss , $f(R,L_{m},T)$ gravity Tangphati:2024war , Weyl geometric gravity Haghani:2023nrm and many others (see Ref. Olmo:2019flu for a review). Quark stars are studied in modified gravity due to the possibility of having larger mass-radius relations in these theories (when compared to general relativity). Recent data of gravitational waves events, like GW 190814 LIGO , points out that one of the components of the coalescence between a black hole and a compact star presents mass $M=2.59^{+0.08}_{-0.09}M_{\odot}$, where $M_{\odot}$ is the solar mass. However, using the MIT bag model and reasonable values for the bag constant (as we will see), general relativity is not able to surpass the $2.4M_{\odot}$ value for a static and noncharged star. On the other hand, a star in the bumblebee gravity could have $M=2.53M_{\odot}$ or even larger values.
This article is structured as follows: In Section II the adopted bumblebee model is presented, the exterior solution, Schwarzschild-like spacetime, is briefly discussed. The analytic interior solution is built. Section III discusses two types of fluid for solving the modified TOV equation. In the incompressible case, approximate expressions for the pressure and the metric inside the star are shown. The second type of fluid under study is the quark matter that the MIT bag model describes. The mass-radius result is exhibited, and we can see the enhancement of this relation due to the bumblebee field. The final comments are in Section IV. In our calculations we adopt geometrized units, i.e., $G=c=1$, where $G$ is the Newtonian constant, and $c$ is speed of light in vacuum. As we present the results, the units in terms of MeV/fm3 are restored.
@@ II.1 The adopted model
$S_{B}$ $=\int d^{4}x\sqrt{-g}\bigg{[}\frac{R}{2\kappa}+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}$
$\ \ \ -V(B^{\mu}B_{\mu}\pm b^{2})+\mathcal{L}_{M}\bigg{]},$ (1)
@@
$=$ $\ \kappa\left[2V^{\prime}B_{\mu}B_{\nu}+B_{\mu}^{\ \alpha}B_{\nu\alpha}-\left(V+\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}\right)g_{\mu\nu}\right]$
$+\xi\bigg{[}\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}$
$+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)$
$-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right)\bigg{]}+\kappa T^{M}_{\mu\nu}.$ (3)
@@ II.2 The exterior solution
$ds^{2}=$ $-\left(1-\frac{2M}{r}\right)dt^{2}+(1+\ell)\left(1-\frac{2M}{r}\right)^{-1}dr^{2}$
$+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (9)
@@ II.3 The interior solution
$\frac{e^{-2\beta}}{r^{2}}\bigg{[}$ $(1+\ell)\left(1+2r\alpha^{\prime}\right)-e^{2\beta}-\ell r^{2}\Big{(}\alpha^{\prime\prime}+\alpha^{\prime 2}$
$-\alpha^{\prime}\beta^{\prime}-\frac{2}{r}\beta^{\prime}\Big{)}\bigg{]}=\kappa p,$ (12)
@@ III.1 Constant density star
$p(r)=\rho_{0}\Bigg{(}\frac{\sqrt{1-\frac{2M}{R}}-\sqrt{1-\frac{2Mr^{2}}{R^{3}}}}{\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}}\Bigg{)}$
$+\ell\rho_{0}\left(\frac{7-\frac{2M}{R}\left(\frac{r^{2}}{R^{2}}+6\right)-7\sqrt{\left(1-\frac{2Mr^{2}}{R^{3}}\right)\left(1-\frac{2M}{R}\right)}}{\left(\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}\right)^{2}}\right),$ (26)
@@
Fig. 1 shows the star pressure (exactly, numerically, and approximately calculated) with the same radius $R=3.2$ km (thus same mass) for different values of the Lorentz-violating parameter $\ell$. As we can see, the bumblebee field decreases the central pressure of the star. This will no longer be the case for the quark matter case in the next section. This effect in the constant-$\rho$ case is very counter-intuitive, because in this type of fluid the pressure decreases once the radius decreases. Using the approximate pressure (26) for $r=0$, it is possible to illustrate the relation between both the central pressure and the parameter $\ell$ in order to see the effect of the bumblebee field on $p_{c}$. In the same graphic (bottom), we see that the approximation (26) is no longer valid for $\ell>10^{-3}$, as we compare it with the pressure graphic (top).
Figure 1: Pressure (top) and central pressure (bottom) for a star with $\rho=\rho_{0}=2$ GeV/fm3 and $R=3.2$ km. In the pressure graphic, exact, numerical, and approximate calculations are compared each other. The bumblebee field decreases the central pressure of the star as $\ell$ increases. However, the approximation breaks for $\ell>10^{-3}$.
@@
$e^{\alpha(r)}=$ $\frac{3}{2}\sqrt{1-\frac{2M}{R}}-\frac{1}{2}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}+\frac{3\ell}{4}\bigg{(}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}$
$-\sqrt{1-\frac{2M}{R}}\bigg{)}.$ (28)
@@
Figure 3: Pressure (top), energy density (middle) and mass function (bottom) with the MIT bag equation of state. The bumblebee field, represented by the Lorentz-violating parameter $\ell$, increases the three quantities. We adopted $\mathcal{B}=60$ MeV/fm3 and $R=11$ km in these graphics.
@@ III.2 Quark matter star
where, as we mentioned, $\mathcal{B}$ is the bag constant. The MIT bag model confines quarks inside a bag, and the bag’s energy, given as $E(r)=\frac{4\pi}{3}r^{3}\mathcal{B}+\frac{C}{r}$, combines a term proportional to the quarks’ kinetic energy and their volume. The bag energy takes place when the energy reaches its minimum, consequently determining the constant $C$. In this context, the bag constant is defined as $\mathcal{B}=E_{0}/(4V_{0})$, with $E_{0}$ representing the energy and $V_{0}$ denoting the volume of the bag. For a neutron, $m\approx E_{0}\approx 1\text{ GeV}$ with a radius of $r_{0}\approx 1\text{ fm}$, resulting in 60 $\mathcal{B}\approx 60\text{ GeV/fm}^{3}$ (or $8\times 10^{-5}$ km-2 in geometrized units) for numerical calculations. This value was used in Refs. Gondek-Rosinska:2008zmv ; Fraga:2013qra ; Arbanil:2016wud . However, given that the MIT bag model is a simple model, it is justifiable to slightly adjust this value. Thus, like Ref. Gondek-Rosinska:2008zmv , we also adopt $\mathcal{B}=40$ MeV/fm3 (equivalent to $5\times 10^{-5}$ km-2 in geometrized units). Unfortunately, for the equation of state (29) analytic solutions are not possible.
@@
Figure 5: Mass-radius relation in the bumblebee context for quark matter described by the MIT bag model with two values for the bag constant $\mathcal{B}$. For large values of the Lorentz-violating parameter $\ell$, the bumblebee gravity could provide larger mass-radius relations than general relativity. The pulsars PSR J0348+0432 Antoniadis:2013pzd $(M=2.01\pm 0.04M_{\odot})$, PSR J0740+6620 Fonseca:2021wxt $(M=2.08\pm 0.07M_{\odot})$, and PSR J0952-0607 Romani:2022jhd $(M=2.35\pm 0.17M_{\odot})$ are described/predicted by general relativity with the bag constant $\mathcal{B}=40$ MeV/fm3. On the other hand, the compact star in the GW 190814 LIGO event, with $M=2.59^{-0.08}_{-0.09}M_{\odot}$, is not. But it could be in the bumblebee context. See Table 1 for a more detailed comparison.
@@
As mentioned in Introduction, the bumblebee model could provide more massive stars than the general relativity context, for the mass-radius relation is also modified due to the bumblebee field and its VEV. As we can see in Fig. 5, large values for the Lorentz breaking parameter $\ell$ increases the mass-radius relation dramatically, except for a threshold (about $\ell=10^{-1}$) from which larger values of $\ell$ do not increase that relation. Therefore, a modified gravitational model like the bumblebee model could fit data for large mass-radius relations in which the general relativity context is not able to fit. In particular, we see in Fig. 5 and Table 1 that for $\mathcal{B}=60$ MeV/fm3 and $\ell=1.9\times 10^{-2}$, the bumblebee model give us a star with $R=2M_{\odot}$. On the other hand, general relativity is not able to reach the $R=2M_{\odot}$ limit. With a small bag constant (as we can also see in Table 1), for example $\mathcal{B}=40$ MeV/fm3, the bumblebee model is able to describe even more massive stars, like the one in the gravitational wave event GW 190814, with mass $M=2.59^{-0.08}_{-0.09}M_{\odot}$. Contrary to general relativity, which does not fit this data ($M=2.45M_{\odot}$), the bumblebee model with $\ell=1.9\times 10^{-2}$ predicts a massive quark star with $M=2.53M_{\odot}$ and $R=13.75$ km.
IV Final remarks

### v3

@@ Abstract
In this work, the interior spacetime of stars is built in a Lorentz symmetry breaking model called bumblebee gravity. Firstly, we calculated the modified Tolman-Oppenheimer-Volkoff equation in this context of modified gravity. Then we show that the bumblebee field, responsible for the symmetry breaking, increases the star mass-radius relation when it assumes its vacuum expectation value. When compared to the general relativity mass-radius relation, a Lorentz symmetry breaking context, like the bumblebee gravity, could provide more massive stars, surpassing the $2.5M_{\odot}$ limit as the interior of the star is described by quark matter with the MIT bag model. Also, we investigate the stability of the solution with the MIT bag equation of state in this context of modified gravity.
@@ I Introduction
In this sense, quark matter in astrophysics has been subject of many papers. In modified theories of gravity, we see works in Einstein-Gauss-Bonnet gravity Tangphati:2021tcy ; Tangphati:2021mvu ; Pretel:2021czp ; Gammon:2023uss ; Pretel:2021gqq , $f(R,L_{m},T)$ gravity Tangphati:2024war , Weyl geometric gravity Haghani:2023nrm and many others (see Ref. Olmo:2019flu for a review). Quark stars are studied in modified gravity due to the possibility of having larger mass-radius relations in these theories (when compared to general relativity). Recent data of gravitational waves events, like GW 190814 LIGO , points out that one of the components of the coalescence between a black hole and a compact star presents mass $M=2.59^{+0.08}_{-0.09}M_{\odot}$, where $M_{\odot}$ is the solar mass. However, using the MIT bag model and reasonable values for the bag constant (as we will see), general relativity is not able to surpass the $2.4M_{\odot}$ value for a static and noncharged star. On the other hand, a star in the bumblebee gravity could have $M=2.53M_{\odot}$ or even larger values. It is worth pointing out that the nature of the compact object in the GW190814 event is still unknown. Besides the quark star hypothesis Tangphati:2024war ; Haghani:2023nrm , there are alternatives that describe such a compact object as a light black hole LIGO ; Nathanail:2021tay , a fast rotating neutron star Zhang:2020zsc ; Most:2020bba ; Biswas:2020xna or even a hybrid star Zhang:2020dfi . Another alternative is to describe the compact object in the GW190814 event as a neutron star in the modified context of the Eddington-inspired Born-Infeld theory Prasetyo:2021kfx . In such a context, a neutron star with $2.6M_{\odot}$ is possible. However, as mentioned, we assume here that the compact object in the GW190814 event could be a quark star in bumblebee gravity.
Recently, three works on bumblebee gravity and stars drawn our attention. The first one Ji:2024aeg builds an interior solution from another exterior solution in bumblebee gravity, namely the spacetime obtained in Bertolami and Páramos Bertolami . In our case, as mentioned before, the exterior solution is the Schwarzschild-like metric of Casana et al. Casana . In this case, due to the fact that different exterior solutions generate different interior solutions, there are no overlapping results. Regarding the two other papers, the authors tried to obtain the interior solution of a star from the Schwarzschild-like metric like us. But the modified TOV is incorrect in either cases. In Ref. Liu , for example, the polytropic fluid studied is not solution of all modified field equations, and in Ref. Panotopoulos:2024jtn even the exterior solution (a vacuum solution) is not solution of the field equations obtained in the mentioned work.
This article is structured as follows: In Section II the adopted bumblebee model is presented, the exterior solution, Schwarzschild-like spacetime, is briefly discussed. The analytic interior solution is built. Section III discusses two types of fluid for solving the modified TOV equation. In the incompressible case, approximate expressions for the pressure and the metric inside the star are shown. The second type of fluid under study is the quark matter that the MIT bag model describes. The mass-radius result is exhibited, and we can see the enhancement of this relation due to the bumblebee field. In Section IV, stability of the solution with the MIT bag equation of state is studied calculating the speed of sound in the fluid, the adiabatic index and the mass-central pressure diagram. The final comments are in Section V. In our calculations we adopt geometrized units, i.e., $G=c=1$, where $G$ is the Newtonian constant, and $c$ is speed of light in vacuum. As we present the results, the units in terms of MeV/fm3 are restored.
@@ II.1 The adopted model
$S_{B}=\int d^{4}x\sqrt{-g}\bigg{[}\frac{R}{2\kappa}+\frac{\xi}{2\kappa}B^{\mu}B^{\nu}R_{\mu\nu}-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-V(B^{\mu}B_{\mu}\pm b^{2})+\mathcal{L}_{M}\bigg{]},$ (1)
@@
$=$ $\ \kappa\left[2V^{\prime}B_{\mu}B_{\nu}+B_{\mu}^{\ \alpha}B_{\nu\alpha}-\left(V+\frac{1}{4}B_{\alpha\beta}B^{\alpha\beta}\right)g_{\mu\nu}\right]+\xi\bigg{[}\frac{1}{2}B^{\alpha}B^{\beta}R_{\alpha\beta}g_{\mu\nu}-B_{\mu}B^{\alpha}R_{\alpha\nu}-B_{\nu}B^{\alpha}R_{\alpha\mu}$
$+\frac{1}{2}\nabla_{\alpha}\nabla_{\mu}\left(B^{\alpha}B_{\nu}\right)+\frac{1}{2}\nabla_{\alpha}\nabla_{\nu}\left(B^{\alpha}B_{\mu}\right)-\frac{1}{2}\nabla^{2}\left(B_{\mu}B_{\nu}\right)-\frac{1}{2}g_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\left(B^{\alpha}B^{\beta}\right)\bigg{]}+\kappa T^{M}_{\mu\nu},$ (3)
@@ II.2 The exterior solution
$ds^{2}=-\left(1-\frac{2M}{r}\right)dt^{2}+(1+\ell)\left(1-\frac{2M}{r}\right)^{-1}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (9)
@@ II.3 The interior solution
$\frac{e^{-2\beta}}{r^{2}}\bigg{[}$ $(1+\ell)\left(1+2r\alpha^{\prime}\right)-e^{2\beta}-\ell r^{2}\Big{(}\alpha^{\prime\prime}+\alpha^{\prime 2}-\alpha^{\prime}\beta^{\prime}-\frac{2}{r}\beta^{\prime}\Big{)}\bigg{]}=\kappa p,$ (12)
@@ III.1 Constant density star
$p(r)=\rho_{0}\left(\frac{\sqrt{1-\frac{2M}{R}}-\sqrt{1-\frac{2Mr^{2}}{R^{3}}}}{\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}}\right)+\ell\rho_{0}\left(\frac{7-\frac{2M}{R}\left(\frac{r^{2}}{R^{2}}+6\right)-7\sqrt{\left(1-\frac{2Mr^{2}}{R^{3}}\right)\left(1-\frac{2M}{R}\right)}}{\left(\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-3\sqrt{1-\frac{2M}{R}}\right)^{2}}\right),$ (26)
@@
Figure 1: Pressure (left) and central pressure (right) for a star with $\rho=\rho_{0}=2$ GeV/fm3 and $R=3.2$ km. In the pressure graphic, exact, numerical, and approximate calculations are compared each other. The bumblebee field decreases the central pressure of the star as $\ell$ increases. However, the approximation breaks for $\ell>10^{-3}$.
@@
Fig. 1 shows the star pressure (exactly, numerically, and approximately calculated) with the same radius $R=3.2$ km (thus same mass) for different values of the Lorentz-violating parameter $\ell$. As we can see, the bumblebee field decreases the central pressure of the star. This will no longer be the case for the quark matter case in the next section. This effect in the constant-$\rho$ case is very counter-intuitive, because in this type of fluid the pressure decreases once the radius decreases. Using the approximate pressure (26) for $r=0$, it is possible to illustrate the relation between both the central pressure and the parameter $\ell$ in order to see the effect of the bumblebee field on $p_{c}$. In the same graphic (right), we see that the approximation (26) is no longer valid for $\ell>10^{-3}$, as we compare it with the pressure graphic (left).
@@
$e^{\alpha(r)}=\frac{3}{2}\sqrt{1-\frac{2M}{R}}-\frac{1}{2}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}+\frac{3\ell}{4}\bigg{(}\sqrt{1-\frac{2Mr^{2}}{R^{3}}}-\sqrt{1-\frac{2M}{R}}\bigg{)}.$ (28)
@@ III.2 Quark matter star
where, as we mentioned, $\mathcal{B}$ is the bag constant. The MIT bag model confines quarks inside a bag, and the bag’s energy, given as $E(r)=\frac{4\pi}{3}r^{3}\mathcal{B}+\frac{C}{r}$, combines a term proportional to the quarks’ kinetic energy and their volume. The bag energy takes place when the energy reaches its minimum, consequently determining the constant $C$. In this context, the bag constant is defined as $\mathcal{B}=E_{0}/(4V_{0})$, with $E_{0}$ representing the energy and $V_{0}$ denoting the volume of the bag. For a neutron, $m\approx E_{0}\approx 1\text{ GeV}$ with a radius of $r_{0}\approx 1\text{ fm}$, resulting in 60 $\mathcal{B}\approx 60\text{ MeV/fm}^{3}$ (or $8\times 10^{-5}$ km-2 in geometrized units) for numerical calculations. This value falls within the range of the bag constant, $\mathcal{B}\sim 57-92$ MeV/fm3, allowing for the reproduction of hadronic masses at low-energy regimes DeTar:1983rw , and has been utilized in Refs. Gondek-Rosinska:2008zmv ; Fraga:2013qra ; Arbanil:2016wud . However, given that the MIT bag model is a simple model, in extreme energy regimes the strong coupling constant for quark interactions is less than that one utilized in DeTar:1983rw . In extreme density regimes, the strong coupling constant is less than in typical conditions where values of $\mathcal{B}$ were derived. In this sense, it is justifiable to slightly adjust the bag constant. Thus, like Ref. Gondek-Rosinska:2008zmv , we also adopt $\mathcal{B}=40$ MeV/fm3 (equivalent to $5\times 10^{-5}$ km-2 in geometrized units). Unfortunately, for the equation of state (29) analytic solutions are not possible.
Figure 3: Pressure (left), energy density (right) and mass function (bottom) with the MIT bag equation of state. The bumblebee field, represented by the Lorentz-violating parameter $\ell$, increases the three quantities. We adopted $\mathcal{B}=60$ MeV/fm3 and $R=11$ km in these graphics.
@@
As mentioned in Introduction, the bumblebee model could provide more massive stars than the general relativity context, for the mass-radius relation is also modified due to the bumblebee field and its VEV. As we can see in Fig. 5, large values for the Lorentz-violating parameter $\ell$ increases the mass-radius relation dramatically, except for a threshold (about $\ell=10^{-1}$) from which larger values of $\ell$ do not increase that relation. Also, it is worth mentioning the “incomplete” curves for large values of $\ell$ in Fig. 5. A given curve in the mass-radius graphic is solution of the TOV with conditions like $p(r)\geq 0$ and $r-2m(r)\geq 0$. For large values of the Lorentz-violating parameter, those conditions are satisfied just for smaller range of parameters. As we will see, this is linked to the stability of the solutions.
Therefore, for appropriate values of the Lorentz-violating parameter $\ell$, a modified gravitational model like the bumblebee model could fit data for large mass-radius relations in which the general relativity context is not able to fit. In particular, we see in Fig. 5 and Table 1 that for $\mathcal{B}=60$ MeV/fm3 and $\ell=1.9\times 10^{-2}$, the bumblebee model give us a star with $R=2M_{\odot}$. On the other hand, general relativity is not able to reach the $R=2M_{\odot}$ limit. With a small bag constant (as we can also see in Table 1), for example $\mathcal{B}=40$ MeV/fm3, the bumblebee model is able to describe even more massive stars, like the one in the gravitational wave event GW 190814, with mass $M=2.59^{-0.08}_{-0.09}M_{\odot}$. Contrary to general relativity, which does not fit this data ($M=2.45M_{\odot}$), the bumblebee model with $\ell=1.9\times 10^{-2}$ predicts a massive quark star with $M=2.53M_{\odot}$ and $R=13.75$ km. This is a virtue of modified theories of gravity Gammon:2023uss ; Tangphati:2024war ; Haghani:2023nrm . Because theories that predict smaller mass-radius relations than general relativity does would be under pressure due to recent data.
Figure 5: On the left, the mass-radius relation in the bumblebee context for quark matter described by the MIT bag model with two values for the bag constant $\mathcal{B}$. For large values of the Lorentz-violating parameter $\ell$, the bumblebee gravity could provide larger mass-radius relations than general relativity. The pulsars PSR J0348+0432 Antoniadis:2013pzd $(M=2.01\pm 0.04M_{\odot})$, PSR J0740+6620 Fonseca:2021wxt $(M=2.08\pm 0.07M_{\odot})$, and PSR J0952-0607 Romani:2022jhd $(M=2.35\pm 0.17M_{\odot})$ are described/predicted by general relativity with the bag constant $\mathcal{B}=40$ MeV/fm3. On the other hand, the compact star in the GW 190814 LIGO event, with $M=2.59^{-0.08}_{-0.09}M_{\odot}$, is not. But it could be in the bumblebee context. See Table 1 for a more detailed comparison. On the right, the $M-\rho_{c}$ curve shows stability just for $\frac{dM}{d\rho_{c}}>0$. The black dots $(\frac{dM}{d\rho_{c}}=0)$ indicate the limit of stable configurations.
@@
IV Stability analysis
In this section, we briefly comment stability of the bumblebee star with quark matter described by the MIT bag equation of state (29). In general relativity, a necessary but not sufficient stability criterion regards the central pressure $\rho_{c}$. It is assumed, even in modified gravity works, that $\frac{dM}{d\rho_{c}}>0$ is a necessary condition to stability in the static case Zeldovich ; Olmo:2019flu . As we can see in Fig. 5, the $M-\rho_{c}$ curve indicates stability intervals for both different values of the bag constant $\mathcal{B}$ and the Lorentz-violating parameter $\ell$. Also we can see that large values of $\ell$ yield small stable ranges. This is in agreement with the mass-radius diagram that shows “incomplete” curves for large values of $\ell$.
It is common to adopt the speed of sound in the fluid, $c_{s}=\sqrt{\frac{dp}{d\rho}}$, to check whether or not causality and stability issues could arise. Inside a stable star, the speed of sound of the fluid must never surpass the speed of light $c$. Using the MIT bag equation of state, it is easy to see that $c_{s}=\frac{c}{\sqrt{3}}$ (where the speed of light in vacuum was restored), that is, one has a constant subluminal speed of sound in the fluid, thus no causality and stability issues are present due to the speed of sound.
Another useful parameter to investigate the star stability is the adiabatic index of perturbations, namely
$\gamma=\left(1+\frac{\rho}{p}\right)\left(\frac{dp}{d\rho}\right)_{S},$ (30)
where $S$ means that the derivative (or the speed of sound in the fluid) is calculated at a constant specific entropy. In the general relativity context, for spherical symmetry, Chandrasekhar Chandrasekhar:1964zz showed using radial perturbations that $\gamma_{cr}=\frac{4}{3}+\frac{19}{42}b$ is the critical value of the adiabatic index for a given stable star, where $b$ is the compactness parameter. That is, stable stars must have $\gamma>\gamma_{cr}$. In bumblebee gravity (and other modified gravity theories) there is no available $\gamma_{cr}$ yet. Thus, it is common to assume that stable stars in these alternative contexts must satisfy at least $\gamma>\frac{4}{3}$, which is the Newtonian limit for the adiabatic index. This is the case here, for using the MIT bag equation of state (29), one has $\gamma=\frac{1}{3}\left(4+\frac{1}{p}\right)>\frac{4}{3}$, according to the positivity of $p$ in Fig. 3, suggesting then stability.
V Final remarks